# New Physics at Tera-$Z$: Precision Renormalised

### v1

@@ 1 Introduction
Specifically, the models considered are all of those which may generate dimension-six SMEFT operators at tree level, as catalogued, with tree-level matching coefficients, in deBlas:2017xtg (the ‘Granada’ catalogue). The novel element considered here is that one-loop renormalisation group (RG) evolution from the new physics scale to the EW scale is accounted for, both in conceptual structural terms and in practical quantitative terms when estimating Tera-$Z$ reach. This RG-evolution brings nearly all models within Tera-$Z$ reach, even if the models don’t contribute to EWPOs at tree level. The RG contribution is IR-calculable and difficult to tune away, since it is proportional to a logarithm of a ratio of scales, unlike finite matching contributions.
@@ 2 EWPOs and their EFT interpretation
The flavour criteria, it turns out, are much more stringent. If we ignore the effects of small Yukawa couplings in the running (i.e. neglecting all but $y_{t}$, a very good approximation), and work in the basis where the up-type Yukawa matrix is diagonal, then the anomalous dimension matrix block-diagonalises according to the flavour structure of the operators Machado:2022ozb . In particular, no operator coefficients with flavour off-diagonal indices can generate the flavour diagonal (or purely bosonic) EWPO coefficients (12). This drastically cuts down the parameter space, and we find that only 168 operator coefficients have the right flavour structure to run into EWPOs (at any loop order). This is only $\sim 7\%$ of the SMEFT parameter space at dimension six.
@@ 3 Categorising states that can escape EWPO constraints
${\omega}_{1}$ $R_{\tau}$ ($R_{\mu}$)
${\omega}_{2}$ $R_{\mu}$
${\omega}_{4}$ $A_{e}$ ($A_{\tau}$)
$\Pi_{1}$ $A_{e}$ ($A_{\tau}$)
@@
$\zeta$ $A_{\tau}$
$\Omega_{1}$ $R_{\mu}$
$\Omega_{2}$ $R_{\mu}$
$\Omega_{4}$ if $(y_{\Omega_{4}})_{33}\neq 0$, all else zero $R_{\mu}$ (-)
$\Upsilon$ $R_{\mu}$
$\Phi$ $R_{\mu}$
@@
$E$ $m_{W}$ ($A_{\tau}$)
$\Delta_{1}$ $A_{e}$ ($A_{\tau}$)
$\Delta_{3}$ $A_{e}$ ($A_{\tau}$)
$\Sigma$ $R_{\mu}$ ($A_{\tau}$)
$\Sigma_{1}$ $A_{e}$ ($A_{\tau}$)
$U$ if $(\lambda_{U})_{3}\neq 0$, all else zero $m_{W}$
$D$ $R_{\mu}$ ($m_{W}$)
@@
${\cal W}$ if $(g_{\mathcal{W}}^{\phi})=0$ $K_{12}(g_{\mathcal{W}}^{l})=0$ if eqns. (27) $A_{e}$ ($A_{\tau}$)
@@
${\cal G}$ if $(g^{u}_{\cal G})_{33}\neq 0$, all else zero $R_{\mu}$
${\cal G}_{1}$ $R_{\mu}$
${\cal H}$ $R_{b}$ ($R_{\mu}$)
${\cal L}_{1}$ $m_{W}$
@@
While it is true that many scalars and vectors may not contribute to the EWPOs at tree level, given the expected precision it is also important to investigate whether any states can escape contributions to EWPOs at loop level. For an extension to the SM to produce zero effect in EWPO operators at one-loop in RGEs, it must additionally produce zeroes in the right-hand sides of all of the following equations Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga : [^3] We omit the RGE for the operator coefficient $\mathcal{C}_{HWB}$ here, as it depends only on coefficients which are not generated at tree level by weakly coupled UV completions deBlas:2017xtg ; Einhorn:2013kja ; Craig:2019wmo , and hence does not imply additional constraints on the couplings of the BSM states.
$\dot{\mathcal{C}}_{HD}$ $=N_{c}\left(\frac{16}{3}Y_{H}Y_{u}g_{1}^{2}-8y_{t}^{2}\right)\mathcal{C}_{Hu}^{33}+\frac{80}{3}g_{1}^{2}Y_{H}^{2}\mathcal{C}_{H\Box},$ (17)
@@
$\dot{\mathcal{C}}_{Hl}^{(3)jj}$ $=\frac{2}{3}g_{2}^{2}\sum_{k}\left(C_{ll}^{jkkj}+N_{c}\mathcal{C}_{lq}^{(3)jjkk}\right),$ (22)
@@
We can compare these to the charge assignments needed such that $\mathcal{B}$ can correspond to the gauge boson of an anomaly-free $U(1)$ gauge symmetry (see e.g. Section 2 of Ref. Allanach:2018vjg ). We find that anomaly-free scenarios can exist, in which the charges of the SM fermions are necessarily non-universal in flavour. For example, a viable anomaly-free charge assignment satisfying (26) is $(g_{\mathcal{B}}^{u})_{11}=(g_{\mathcal{B}}^{d})_{11}=(g_{\mathcal{B}}^{e})_{11}=-(g_{\mathcal{B}}^{u})_{22}=-(g_{\mathcal{B}}^{d})_{22}=-(g_{\mathcal{B}}^{e})_{22}$, $(g_{\mathcal{B}}^{u})_{33}=(g_{\mathcal{B}}^{d})_{33}=(g_{\mathcal{B}}^{e})_{33}=0$. It turns out that it is not possible to obtain anomaly-free charge assignments satisfying (26) which are universal in the first two generations, and therefore any misalignment between the flavour and gauge bases will lead to tree level flavour changing neutral currents (FCNCs) among the light generations of quarks and/or leptons.
@@ 4 Probing one-particle extensions of the SM
As discussed in the previous section, the majority of one-particle extensions of the SM give rise, either at tree or at loop level, to effects in EWPOs. We now turn to quantitatively analysing the different options one by one, i.e. we consider one new state at a time, from the list of Ref. deBlas:2017xtg . In each case we study the projected constraints from a Tera-$Z$ run at FCC-ee, taking the sensitivities from Refs. DeBlas:2019qco ; Bernardi:2022hny . We adopt two benchmark scenarios for the flavour structure of the couplings to the SM fermions:
@@
In the conventions of Ref. deBlas:2017xtg we set all (non-gauge) dimensionless couplings to unity, in natural units, for the sake of simplicity of presentation. For all trilinear scalar couplings we choose the scale $\kappa=5$ TeV, in order to avoid the additional suppression in $\kappa/M$ contributions to the dimension-six coefficients. Note that perturbativity is preserved in those cases, given the values of the masses we obtain. We do not include any interactions involving the new states at mass dimension five or above. We find that this would not affect our qualitative conclusions, since no additional SMEFT operators which contribute at tree or one loop level to EWPOs are generated through these couplings, for any of the states.
@@
We calculate the pattern of Wilson coefficients at the EW scale by running the full set of tree-level operators from the matching scale down to the EW scale, thus capturing the leading-log IR-calculable contributions to the EWPOs. Note that for consistency this is done also in the cases for which constributions to EWPOs arise already at tree level. [^6] We find that inclusion of RG effects in those cases leads to a slight modification of the bounds, with a few exceptions for which one-loop effects dominate (see detailed discussion below). Running is calculated at the first leading-log level and is, for indicative purposes, run from a scale of $2$ TeV down to the EW scale using DsixTools Celis:2017hod ; Fuentes-Martin:2020zaz . This scale has been chosen to roughly correspond to the average bounds that we find on states contributing to EWPOs via RGEs only.
@@
This RG evolution leads, in almost all cases, to non-vanishing contributions to the operators responsible for EWPOs. Unless otherwise stated the one-loop matching contributions from within the UV theory are not included. We opt for this choice for a number of reasons. The first is that we wish to answer the question of whether a UV completion can generate dimension-six SMEFT contributions without contributing to EWPOs at the one-loop level. Since we can, in almost all cases, answer this question in the affirmative already with IR-calculable effects it is not necessary to to appeal further to one-loop matching terms. The second reason is that these contributions are very model-dependent, sensitive to a much wider range of BSM couplings than the tree-level contributions, hence assigning too much importance to their numerical value when working in the couplings $\to 1$ assumption would be taking the models too seriously at the quantitative level, especially when it is qualitative lessons we aim to extract. Note that for one-loop matching one opens up sensitivity to all couplings which preserve a discrete $\mathcal{Z}_{2}$ symmetry acting on the new states, unlike for the tree-level terms. Finally, we are not aware of any automated package capable of calculating the full set of matching coefficients in the Warsaw basis for all of the models considered in this work  [^7] Although automated software packages for one-loop matching such as Matchete Fuentes-Martin:2022jrf and Matchmakereft Carmona:2021xtq exist, they are currently not able to treat the integration of vector states consistently, which is why we refrain from using them in this work., without which such an extensive programme of calculations would be prohibitively onerous considering the minor quantitative and negligible qualitative impact on the results. The only exception is for the models $\Theta_{1}$ and $\Theta_{3}$, for which the leading contributions to EWPOs arise at the one-loop matching level and are included here, for the sake of demonstrating that they, too, do not evade EWPOs at one-loop.
@@
Figure 3: Projected bounds on the masses of new scalar fields. The vertical dashed lines separate fields which contribute to EWPOs at tree level (left), via one-loop RG evolution (middle), and via one-loop matching (right). The green and blue bars correspond to different assumptions for the coupling to SM fermions, as described near the start of Sec. 4, while red bars are used when the state does not have couplings to fermions. Fields indicated with a (∗) correspond to cases where the tree-level contribution has been set to zero by forbidding a specific coupling (see Table 3). Hatched bars correspond to pure tree-level limits, without RG running.
@@
Figure 4: Projected bounds on the masses of new vector fields. The vertical dashed line separates fields which contribute to EWPOs at tree level (left) and via one-loop RG evolution (right). The green and blue bars correspond to different assumptions for the coupling to SM fermions, as described near the start of Sec. 4, while red bars are used when the state does not have couplings to fermions. Fields indicated with a (∗) correspond to cases where the tree-level contribution has been set to zero by forbidding a specific coupling (see Table 5). Hatched bars correspond to pure tree-level limits, without RG running.
@@
Figure 5: Projected bounds on the masses of new fermion fields. The vertical dashed line separates fields which contribute to EWPOs at tree level (left) and via one-loop RG evolution (right). The green and blue bars correspond to different assumptions for the coupling to SM fermions, as described near the start of Sec. 4, while the red bar for $Q_{1}^{(}*)$ corresponds to the exceptional case where $Q_{1}$ couples only to right-handed top quarks, as noted in Table 4. Fields indicated with a (∗) correspond to cases where the tree-level contribution has been set to zero by forbidding a specific coupling (see Table 4). Hatched bars correspond to pure tree-level limits, without RG running.
@@ Acknowledgements.
We are grateful to Jorge de Blas, Joe Davighi, Christophe Grojean, Gino Isidori, Ben Stefanek and Tevong You for discussions. SR is supported by UKRI Stephen Hawking Fellowship EP/W005433/1, and partially supported by STFC grant ST/X000605/1. The work of LA is funded by the Swiss National Science Foundation (SNF) under contract 200020_204428.

### v2

@@ 1 Introduction
Specifically, the models considered are all of those which may generate dimension-six SMEFT operators at tree level, as catalogued, with tree-level matching coefficients, in deBlas:2017xtg (the ‘Granada’ catalogue). Future limits on these models, including from FCC-ee, have been studied without running in the recent global fits of Ref. Celada:2024mcf . The novel element considered here is that one-loop renormalisation group (RG) evolution from the new physics scale to the EW scale is accounted for, both in conceptual structural terms and in practical quantitative terms when estimating Tera-$Z$ reach. This RG-evolution brings nearly all models within Tera-$Z$ reach, even if the models don’t contribute to EWPOs at tree level. The RG contribution is IR-calculable and difficult to tune away, since it is proportional to a logarithm of a ratio of scales, unlike finite matching contributions.
@@ 2 EWPOs and their EFT interpretation
The flavour criteria, it turns out, are much more stringent. If we ignore the effects of small Yukawa couplings in the running (i.e. neglecting all but $y_{t}$, a very good approximation), and work in the basis where the up-type Yukawa matrix is diagonal, then the anomalous dimension matrix block-diagonalises according to the flavour structure of the operators Machado:2022ozb . In particular, no operator coefficients with flavour off-diagonal indices can generate the flavour diagonal (or purely bosonic) EWPO coefficients (12). This drastically cuts down the parameter space, and we find that only 168 operator coefficients have the right flavour structure to run into EWPOs. This counting is true at any loop order, since it is based on the flavour symmetry structure of SMEFT operators and SM interactions. [^3] To be specific, this counting includes: all CP-even flavour-singlet operators, all CP-even flavour-diagonal 4-fermion operators including a flavour-singlet current or a current with the flavour charges of $y_{t}^{2}$, the CP-even coefficients of $C_{uW}^{33}$, $C_{uB}^{33}$, $C_{uG}^{33}$, $C_{uH}^{33}$, and two extra flavour-diagonal components of $C_{ll}$. For details of the flavour charges of the SMEFT operators, see Machado:2022ozb . This is only $\sim 7\%$ of the SMEFT parameter space at dimension six.
@@ 3 Categorising states that can escape EWPO constraints
${\omega}_{1}$ $\Gamma_{Z}$
${\omega}_{2}$ $\Gamma_{Z}$
${\omega}_{4}$ $A_{e}$ ($R_{\tau}$)
$\Pi_{1}$ $A_{e}$ ($R_{\tau}$)
@@
$\zeta$ $\Gamma_{Z}$ ($R_{\tau}$)
$\Omega_{1}$ $\Gamma_{Z}$
$\Omega_{2}$ $\Gamma_{Z}$
$\Omega_{4}$ if $(y_{\Omega_{4}})_{33}\neq 0$, all else zero $\Gamma_{Z}$ (-)
$\Upsilon$ $\Gamma_{Z}$
$\Phi$ $\Gamma_{Z}$
@@
$E$ $\Gamma_{Z}$ ($R_{\tau}$)
$\Delta_{1}$ $A_{e}$ ($R_{\tau}$)
$\Delta_{3}$ $A_{e}$ ($R_{\tau}$)
$\Sigma$ $\Gamma_{Z}$ ($R_{\tau}$)
$\Sigma_{1}$ $A_{e}$ ($\Gamma_{Z}$)
$U$ if $(\lambda_{U})_{3}\neq 0$, all else zero $\Gamma_{Z}$ ($m_{W}$)
$D$ $\Gamma_{Z}$
@@
${\cal W}$ if $(g_{\mathcal{W}}^{\phi})=0$ $K_{12}(g_{\mathcal{W}}^{l})=0$ if eqns. (27) $A_{e}$ ($\Gamma_{Z}$)
@@
${\cal G}$ if $(g^{u}_{\cal G})_{33}\neq 0$, all else zero $\Gamma_{Z}$
${\cal G}_{1}$ $\Gamma_{Z}$
${\cal H}$ $\Gamma_{Z}$
${\cal L}_{1}$ $A_{e}$
@@
While it is true that many scalars and vectors may not contribute to the EWPOs at tree level, given the expected precision it is also important to investigate whether any states can escape contributions to EWPOs at loop level. For an extension to the SM to produce zero effect in EWPO operators at one-loop in RGEs, it must additionally produce zeroes in the right-hand sides of all of the following equations Jenkins:2013zja ; Jenkins:2013wua ; Alonso:2013hga : [^4] We omit the RGE for the operator coefficient $\mathcal{C}_{HWB}$ here, as it depends only on coefficients which are not generated at tree level by weakly coupled UV completions deBlas:2017xtg ; Einhorn:2013kja ; Craig:2019wmo , and hence does not imply additional constraints on the couplings of the BSM states. [^5] We include a term dependent on $\mathcal{C}_{Hq}^{(1)33}$ in Eqn. (17). On its own, this coefficient induces $Z\to\bar{b}b$ and hence would already contribute at tree level to EWPOs. However, in combination with an equal and opposite contribution to $\mathcal{C}_{Hq}^{(3)33}$ (which does not enter the $\mathcal{C}_{HD}$ RGE) the $Z\bar{b}b$ coupling exactly cancels. So we include this term to cover this possibility.
$\dot{\mathcal{C}}_{HD}$ $=N_{c}\left(\frac{16}{3}Y_{H}Y_{u}g_{1}^{2}-8y_{t}^{2}\right)\mathcal{C}_{Hu}^{33}+N_{c}\left(\frac{32}{3}Y_{H}Y_{q}g_{1}^{2}+8y_{t}^{2}\right)\mathcal{C}_{Hq}^{(1)33}$
$+\frac{80}{3}g_{1}^{2}Y_{H}^{2}\mathcal{C}_{H\Box},$ (17)
@@
$\dot{\mathcal{C}}_{Hl}^{(3)jj}$ $=\frac{2}{3}g_{2}^{2}\sum_{k}\left(\mathcal{C}_{ll}^{jkkj}+N_{c}\mathcal{C}_{lq}^{(3)jjkk}\right),$ (22)
@@
We can compare these to the charge assignments needed such that $\mathcal{B}$ can correspond to the gauge boson of an anomaly-free $U(1)$ gauge symmetry (see e.g. Section 2 of Ref. Allanach:2018vjg ). We find that anomaly-free scenarios can exist, in which the charges of the SM fermions are necessarily non-universal in flavour. For example, a viable anomaly-free charge assignment satisfying (26) is $(g_{\mathcal{B}}^{u})_{11}=(g_{\mathcal{B}}^{d})_{11}=(g_{\mathcal{B}}^{e})_{11}=-(g_{\mathcal{B}}^{u})_{22}=-(g_{\mathcal{B}}^{d})_{22}=-(g_{\mathcal{B}}^{e})_{22}$, $(g_{\mathcal{B}}^{u})_{33}=(g_{\mathcal{B}}^{d})_{33}=(g_{\mathcal{B}}^{e})_{33}=0$. It turns out that it is not possible to obtain anomaly-free charge assignments satisfying (26) which are universal in the first two generations, and therefore any misalignment between the flavour and gauge bases will lead to tree level flavour changing neutral currents (FCNCs) among the light generations of quarks and/or leptons. [^7] We thank Joe Davighi for providing a proof of this fact, and the viable anomaly-free charges we quote here.
@@ 4 Probing one-particle extensions of the SM
As discussed in the previous section, the majority of one-particle extensions of the SM give rise, either at tree or at loop level, to effects in EWPOs. We now turn to quantitatively analysing the different options one by one, i.e. we consider one new state at a time, from the list of Ref. deBlas:2017xtg . In each case we study the projected constraints from a Tera-$Z$ run at FCC-ee, taking the sensitivities from Refs. deBlas:2022ofj . We adopt three benchmark scenarios for the flavour structure of the couplings to the SM fermions:
@@
  • (c)
    Antisymmetric couplings. Some of the scalar extensions considered are forced by gauge symmetry to have an antisymmetric coupling matrix in flavour space. For this reason, choosing either of the two options above would not be consistent, and in particular the third-generation only scenario is not defined. In order to show that these cases still retain sensitivity to EWPOs, we choose instead to switch on all couplings, such that flavour-conserving operators are generated by the product of two flavour-violating couplings. This is the case for $S_{1}$ and $\omega_{2}$ in Fig. 3. Notice that there are other instances of such antisymmetric couplings, namely the di-quark couplings of $\omega_{4}$, $\zeta$, and $\Omega_{1}$. Since in those cases also other couplings (with no particular symmetry structure) are present, we set the antisymmetric ones to zero [^9] Notice that this choice doesn’t change the sensitivity to EWPOs, cf. Tab. 3.
@@
In the conventions of Ref. deBlas:2017xtg we set all (non-gauge) dimensionless couplings to unity, in natural units, for the sake of simplicity of presentation. For all trilinear scalar couplings we choose the scale $\kappa=5$ TeV, in order to avoid the additional suppression in $\kappa/M$ contributions to the dimension-six coefficients. Note that perturbativity is preserved in those cases, given the values of the masses we obtain. We do not include any interactions involving the new states at mass dimension five or above. We find that this would not affect our qualitative conclusions, since these couplings never change whether a state contributes at tree or one loop level to EWPOs.
@@
We calculate the pattern of Wilson coefficients at the EW scale by running the full set of tree-level operators from the matching scale down to the EW scale, thus capturing the leading-log IR-calculable contributions to the EWPOs. Note that for consistency this is done also in the cases for which contributions to EWPOs arise already at tree level. [^10] In these cases we also show pure tree-level bounds without RG effects, as hatched bars on the constraint plots. We find that inclusion of RG effects in these cases leads to a slight modification of the bounds, with a few exceptions for which one-loop effects dominate (see detailed discussion below). Running is calculated at the first leading-log level and is, for indicative purposes, run from a scale of $2$ TeV down to the EW scale using DsixTools Celis:2017hod ; Fuentes-Martin:2020zaz . This scale has been chosen to roughly correspond to the average bounds that we find on states contributing to EWPOs via RGEs only.
@@
This RG evolution leads, in almost all cases, to non-vanishing contributions to the operators responsible for EWPOs. Unless otherwise stated the one-loop matching contributions from within the UV theory are not included. We opt for this choice for a number of reasons. The first is that we wish to answer the question of whether a UV completion can generate dimension-six SMEFT contributions without contributing to EWPOs at the one-loop level. Since we can, in almost all cases, answer this question in the affirmative already with IR-calculable effects it is not necessary to appeal further to one-loop matching terms. The second reason is that these contributions are very model-dependent, sensitive to a much wider range of BSM couplings than the tree-level contributions, hence assigning too much importance to their numerical value when working in the couplings $\to 1$ assumption would be taking the models too seriously at the quantitative level, especially when it is qualitative lessons we aim to extract. Note that for one-loop matching one opens up sensitivity to all couplings which preserve a discrete $\mathcal{Z}_{2}$ symmetry acting on the new states, unlike for the tree-level terms. Finally, we are not aware of any automated package capable of calculating the full set of matching coefficients in the Warsaw basis for all of the models considered in this work  [^11] Although automated software packages for one-loop matching such as Matchete Fuentes-Martin:2022jrf and Matchmakereft Carmona:2021xtq exist, they are currently not able to treat the integration of vector states consistently, which is why we refrain from using them in this work., without which such an extensive programme of calculations would be prohibitively onerous considering the minor quantitative and negligible qualitative impact on the results. The only exception is for the models $\Theta_{1}$ and $\Theta_{3}$, for which the leading contributions to EWPOs arise at the one-loop matching level and are included here, for the sake of demonstrating that they, too, do not evade EWPOs at one-loop.
@@
Figure 3: Projected bounds (95% CL) on the masses of new scalar fields. The vertical dashed lines separate fields which contribute to EWPOs at tree level (left), via one-loop RG evolution (middle), and via one-loop matching (right). The green and blue bars correspond to different assumptions for the coupling to SM fermions, as described near the start of Sec. 4, while red bars are used when the state does not have couplings to fermions. Fields indicated with a (∗) correspond to cases where the tree-level contribution has been set to zero by forbidding a specific coupling (see Table 3). Hatched bars correspond to pure tree-level limits, without RG running.
@@
Figure 4: Projected bounds (95% CL) on the masses of new vector fields. The vertical dashed line separates fields which contribute to EWPOs at tree level (left) and via one-loop RG evolution (right). The green and blue bars correspond to different assumptions for the coupling to SM fermions, as described near the start of Sec. 4, while red bars are used when the state does not have couplings to fermions. Fields indicated with a (∗) correspond to cases where the tree-level contribution has been set to zero by forbidding a specific coupling (see Table 5). Hatched bars correspond to pure tree-level limits, without RG running.
@@
Figure 5: Projected bounds (95% CL) on the masses of new fermion fields. The vertical dashed line separates fields which contribute to EWPOs at tree level (left) and via one-loop RG evolution (right). The green and blue bars correspond to different assumptions for the coupling to SM fermions, as described near the start of Sec. 4, while the red bar for $Q_{1}^{(}*)$ corresponds to the exceptional case where $Q_{1}$ couples only to right-handed top quarks, as noted in Table 4. Fields indicated with a (∗) correspond to cases where the tree-level contribution has been set to zero by forbidding a specific coupling (see Table 4). Hatched bars correspond to pure tree-level limits, without RG running.
@@ Acknowledgements.
We are grateful to Jorge de Blas, Joe Davighi, Admir Greljo, Christophe Grojean, Gino Isidori, Ben Stefanek and Tevong You for discussions. SR is supported by UKRI Stephen Hawking Fellowship EP/W005433/1, and partially supported by STFC grant ST/X000605/1. The work of LA is funded by the Swiss National Science Foundation (SNF) under contract 200020_204428.